OpenStudy (anonymous):

Find all solutions of the equation x^2-4x+7=0 and express them in the form : a+bi

OpenStudy (anonymous):

\[x = 2 + i \sqrt{3}\]

OpenStudy (anonymous):

\[x = 2 - i \sqrt{3}\]

OpenStudy (anonymous):

This is a quadratic so you have two answers.